# Distributionally and Adversarially Robust Logistic Regression via Intersecting Wasserstein Balls

### v1

@@ Abstract
Empirical risk minimization often fails to provide robustness against adversarial attacks in test data, causing poor out-of-sample performance. Adversarially robust optimization (ARO) has thus emerged as the de facto standard for obtaining models that hedge against such attacks. However, while these models are robust against adversarial attacks, they tend to suffer severely from overfitting. To address this issue for logistic regression, we study the Wasserstein distributionally robust (DR) counterpart of ARO and show that this problem admits a tractable reformulation. Furthermore, we develop a framework to reduce the conservatism of this problem by utilizing an auxiliary dataset (e.g., synthetic, external, or out-of-domain data), whenever available, with instances independently sampled from a nonidentical but related ground truth. In particular, we intersect the ambiguity set of the DR problem with another Wasserstein ambiguity set that is built using the auxiliary dataset. We analyze the properties of the underlying optimization problem, develop efficient solution algorithms, and demonstrate that the proposed method consistently outperforms benchmark approaches on real-world datasets.
@@ 1 Introduction
Figure 1: ARO optimizes the expected adversarial loss over the empirical distribution $\mathbb{P}_{N}$ that is constructed from $N$ i.i.d. samples of the (unknown) true data-generating distribution $\mathbb{P}^{0}$. Replacing $\mathbb{P}_{N}$ with the worst-case distribution in a ball $\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})$ gives us its DR counterpart. To reduce the size of this ball while ensuring $\mathbb{P}^{0}$ is still included with high confidence, we intersect it with another ball $\mathfrak{B}_{\widehat{\varepsilon}}(\widehat{P}_{\widehat{N}})$. The latter ball is built around an empirical distribution ^$\widehat{\mathbb{P}}_{N}$ that is constructed from $\widehat{N}$ i.i.d. samples of some auxiliary data distribution $\widehat{\mathbb{P}}$ (cf. Section 5). The intersection includes $\mathbb{P}^{0}$ if $\varepsilon\geq(1)$ and $\widehat{\varepsilon}\geq(2)+(3)$ (cf. Section 6). Recent works using auxiliary data in ARO propose optimizing the expected adversarial loss over a mixture $\mathbb{Q}_{\mathrm{mix}}$ of $\mathbb{P}_{N}$ and $\widehat{P}_{\widehat{N}}$; we show that this distribution also resides in $\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}({\widehat{\mathbb{P}}}_{\widehat{N}})$ under some conditions (Proposition 5.8).
Supervised learning traditionally involves access to a training dataset whose instances are assumed to be independently sampled from a true data-generating distribution [12, 32]. Optimizing an expected loss for the empirical distribution constructed from such a training set, also known as empirical risk minimization (ERM), enjoys several desirable properties in relatively generic settings, including convergence to the true risk minimization problem as the number of training samples increases [72, Chapter 2]. In practice, however, data is finite, and ERM suffers from the “optimism bias” that is also known as overfitting [43] or the optimizer’s curse [20, 63], which causes deteriorated out-of-sample performance. A popular paradigm to prevent this phenomenon is distributionally robust optimization (DRO) [19] which optimizes the expected loss for the worst-case distribution that resides in an ambiguity set constructed from the empirical distribution.
Another actively studied real-world challenge causing poor out-of-sample performance of ERM is adversarial attacks, where an adversary perturbs the observed features in the testing or deployment phase [67, 28]. For neural networks, the paradigm of adversarial training (AT) [38] is therefore designed to provide adversarial robustness by simulating the attacks during the training stage. Many successful variants of AT specialized to different domains, losses, and attacks have been proposed in the literature to achieve adversarial robustness without significantly deteriorating the performance on training sets [56, 81, 45, 27]. While some works (e.g., [14, 70]) examine adversarial robustness guarantees of various training algorithms, there is a recent stream of work (e.g., [6, 76]) that studies properties of optimal solutions to the adversarially robust optimization (ARO) problems where one optimizes the empirical risk subject to worst-case adversarial attacks.
Recently, it has been observed that adversarially robust (AR) models may suffer from severe overfitting (robust overfitting, [49, 79, 37]), that is, AR models are not DR. Indeed, it is observed that robust overfitting is even more severe than traditional overfitting [51]. While some works address robust overfitting of AT through algorithmic adjustments [16, 36], a recent study [6, Thm 3.2] proves that robust overfitting is more severe than traditional overfitting via DRO theory. The authors of the latter work thus propose the simultaneous adoption of DR and AR.
In this paper, we adopt a Wasserstein DRO approach to address robust overfitting in the $\ell_{p}$-attack setting [18] for logistic regression. We study both the traditional setting with an empirical dataset and an extension that incorporates an auxiliary dataset whose instances are sampled from a nonidentical but related distribution. Examples of auxiliary data include synthetic data generated from a generative model (e.g., releasing portions of data under privacy constraints), data in the presence of distributional shifts (e.g., different time period/geographic region), noisy data (e.g., measurement errors), or out-of-domain data (e.g., different source); any distribution is applicable as long as the Wasserstein distance between its underlying data-generating distribution and the true data-generating distribution is known or can be estimated (formal setup in Section 5). We propose a distributionally and adversarially robust model, constructing its ambiguity set from empirical and auxiliary datasets. Specifically, we first develop a Wasserstein DR counterpart of ARO without auxiliary data, which already improves the benchmark ARO methods. Our primary contribution, however, is intersecting this empirical Wasserstein ambiguity set (ball) with an additional ball formed around the auxiliary data. This method mitigates conservatism in DRO by refining its ambiguity set. We analyze the statistical properties and complexities attributed to this problem, and develop efficient approximation algorithms. Figure 1 illustrates the idea and Appendix A contains notation. Our contributions are:
  • We show that ARO for logistic loss is equivalent to the ERM of a new loss function, which is convex and Lipschitz, allowing us to use recent Lipschitz DRO theory (cf. Section 4).
  • We thus formulate distributionally and adversarially robust logistic regression (LR) and provide an exact tractable convex optimization reformulation (cf. Section 4).
  • We utilize auxiliary data to reduce the conservatism of the aforementioned DRO problem in Section 5 (cf. Figure 1). We prove that the resulting optimization problem is NP-hard and develop a tractable approximation.
  • We prove that Wasserstein finite sample guarantees are inherited by our optimization models and discuss how to set the radii of the Wasserstein balls (cf. Section 6).
  • Experiments on UCI datasets and MNIST/EMNIST datasets demonstrate that our approach achieves better out-of-sample performance than benchmark algorithms with and without adversarial attacks, and scales graciously in practical settings (cf. Section 7).
@@ 2 Related work
Despite the difference in motivation from ours, auxiliary data appears in the ARO literature. In particular, it is shown that additional unlabeled data sampled from the same [15, 75] or different [21] data-generating distributions could improve adversarial robustness. [54] shows adversarial robustness guarantees can be certified even when AT is done on a synthetic dataset if its generator’s distance to the true distribution can be quantified. [30, 76] propose optimizing a weighted combination of ARO over empirical and synthetic datasets. We show that the latter approach is generalized by our model (cf. Proposition 5.8).
DRO-ARO interactions
@@
In our work, we solve ARO for the worst-case data distribution residing in a type-1 Wasserstein ball around the empirical distribution, since the type-1 Wasserstein metric is arguably the most common choice in machine learning (ML) with Lipschitz losses [59, 26]. In the literature, it is shown that the standard (non-DR) ARO is equivalent to the DRO of the original loss function with a type-$\infty$ Wasserstein metric [65, 33, 50] (or a Lévy-Prokhorov metric [7]). Hence, our DR ARO approach can be interpreted as optimizing the logistic loss over the worst-case distribution whose $1$-Wasserstein distance is bounded by a pre-specified radius from at least one distribution that resides in an $\infty$-Wasserstein ball around the empirical distribution. Conversely, [62] discusses that while DRO over Wasserstein balls is intractable for generic functions (e.g., neural networks), its Lagrange relaxation resembles ARO and thus AT yields a certain degree of (relaxed) distributional robustness; this introduces a DRO perspective to AT algorithms [74, 13, 47]. However, to the best of our knowledge, there have not been works optimizing a pre-specified level of Wasserstein distributional robustness (that hedges against overfitting, [34]) and adversarial robustness (that hedges against adversarial attacks, [28]) simultaneously. To our knowledge, the only work that considers the DR counterpart of ARO is [6] where the distributional ambiguity is modeled with $\varphi$-divergences and the prediction model is a neural network.
Intersecting ambiguity sets in DRO
@@
Recent work started to explore the intersection of ambiguity sets for different contexts [2] or different metrics [82]. Our idea of intersecting Wasserstein balls is inspired by the “Surround, then Intersect” strategy [68, §5.2] to train linear regression under sequential domain adaptation in a non-adversarial setting (see [57] and [64] for robustness in domain adaptation/transfer learning). The aforementioned work focuses on a case where the loss function is the squared loss, and the metric is a variant of the Wasserstein metric developed for the first and second distributional moments.
Logistic Loss in DRO and ARO
@@
Our choice of LR aligns with the current directions and open questions in the relevant literature. In the ARO literature, there are recent theory developments on understanding the effect of auxiliary data (e.g.[76]) for squared and logistic loss functions. In the DRO literature, even in the absence of adversarial attacks, the aforementioned work [68] on the intersection of Wasserstein ambiguity sets is restricted to linear regression. The authors show that this problem admits a tractable convex optimization reformulation, and the proof relies on the properties of the squared loss. We contribute to the DRO literature for adversarial and non-adversarial settings because we show that such a problem would be NP-hard for the logistic loss (cf., Proposition 5.3), and develop specialized approximation techniques. Our problem recovers DR LR [60, 55] as a special case in the absence of adversarial attacks and auxiliary data. The theoretical challenges posed by the logistic loss have been a significant focus in DRO literature, with extensions such as DR LR [60] to DRO Lipschitz ML [59] and mixed-feature DR LR [55] to mixed-feature DR Lipschitz ML [3].
3 Problem setting and preliminaries
@@
We consider a binary classification problem where an instance is modeled as $(\bm{x},y)\in\Xi:=\mathbb{R}^{n}\times\{-1,+1\}$ and the labels depend on the features probabilistically with $\mathrm{Prob}[y\mid\bm{x}]\;=\;[1+\exp(-y\cdot\bm{\beta}^{\top}\bm{x})]^{-1}$, for some $\bm{\beta}\in\mathbb{R}^{n}$; its associated loss is the logloss $\ell_{\bm{\beta}}(\bm{x},y):=\log(1+\exp{(-y\cdot\bm{\beta}^{\top}\bm{x})})$.
Distributional ambiguity and Wasserstein balls
Let $\mathcal{P}(\Xi)$ denote the set of probability distributions on $\Xi$. We model distributional ambiguity via the Wasserstein (Earth mover’s) distance.
Definition 3.1 (Feature-label metric).
The distance $d(\bm{\xi},\bm{\xi^{\prime}})$ between two instances $\bm{\xi}=(\bm{x},y)\in\Xi$ and $\bm{\xi^{\prime}}=(\bm{x^{\prime}},y^{\prime})\in\Xi$ is
$d(\bm{\xi},\bm{\xi^{\prime}})=\lVert\bm{x}-\bm{x^{\prime}}\rVert_{q}+\kappa\cdot\mathds{1}\small{[y\neq y^{\prime}]},$
where $\kappa>0$ controls the label weight and $q>0$ specifies a rational norm on $\mathbb{R}^{n}$.
Definition 3.2.
The type-1 Wasserstein distance between distributions $\mathbb{Q}\in\mathcal{P}(\Xi)$ and $\mathbb{Q}^{\prime}\in\mathcal{P}(\Xi)$, with ground metric $d(\bm{\xi},\bm{\xi^{\prime}})$ on $\Xi$, is defined as
$\mathrm{W}(\mathbb{Q},\mathbb{Q}^{\prime})=\underset{\Pi\in\mathcal{C}(\mathbb{Q},\mathbb{Q}^{\prime})}{\inf}\left\{\int_{\Xi\times\Xi}d(\bm{\xi},\bm{\xi^{\prime}})\Pi(\mathop{}\!\mathrm{d}\bm{\xi},\mathop{}\!\mathrm{d}\bm{\xi^{\prime}})\right\}$
where $\mathcal{C}(\mathbb{Q},\mathbb{Q}^{\prime}):=\{\Pi\in\mathcal{P}(\Xi\times\Xi):\Pi(\mathop{}\!\mathrm{d}\bm{\xi},\Xi)=\mathbb{Q}(\mathop{}\!\mathrm{d}\bm{\xi}),\Pi(\Xi,\mathop{}\!\mathrm{d}\bm{\xi^{\prime}})=\mathbb{Q}^{\prime}(\mathop{}\!\mathrm{d}\bm{\xi^{\prime}})\}$.
For $\varepsilon>0$, the Wasserstein ball around $\mathbb{P}\in\mathcal{P}(\Xi)$ is defined $\mathfrak{B}_{\varepsilon}(\mathbb{P}):=\{\mathbb{Q}\in\mathcal{P}(\Xi):\mathrm{W}(\mathbb{Q},\mathbb{P})\leq\varepsilon\}$. We next review several training paradigms, see Table 1.
Table 1: Comparison of the risks taken under various training paradigms.
4 6 8
Training risk $\mathbb{E}_{\mathbb{P}_{N}}[\ell_{\bm{\beta}}(\bm{x},y)]$ $\underset{\mathbb{Q}\in\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})}{\sup}\;\mathbb{E}_{\mathbb{Q}}[\ell_{\bm{\beta}}(\bm{x},y)]$ $\mathbb{E}_{\mathbb{P}_{N}}\big{[}\underset{\bm{z}:\lVert\bm{z}\rVert_{p}\leq\alpha}{\sup}\{\ell_{\bm{\beta}}(\bm{x}+\bm{z},y)\}\big{]}$
True risk $\mathbb{E}_{\mathbb{P}^{0}}[\ell_{\bm{\beta}}(\bm{x},y)]$ $\mathbb{E}_{\mathbb{P}^{0}}[\ell_{\bm{\beta}}(\bm{x},y)]$ $\mathbb{E}_{\mathbb{P}^{0}}\big{[}\underset{\bm{z}:\lVert\bm{z}\rVert_{p}\leq\alpha}{\sup}\{\ell_{\bm{\beta}}(\bm{x}+\bm{z},y)\}\big{]}$
Empirical Risk Minimization
@@
Let $\mathbb{P}^{0}$ denote the true data-generating distribution. Ideally, one wants to minimize the expected loss over $\mathbb{P}^{0}$, or more precisely
$\begin{array}[]{cl}\displaystyle\underset{\bm{\beta}\in\mathbb{R}^{n}}{\inf}&\displaystyle\mathbb{E}_{\mathbb{P}^{0}}[\ell_{\bm{\beta}}(\bm{x},y)].\end{array}$ (2)
@@
In practice, $\mathbb{P}^{0}$ is hardly ever known, and one thus resorts to the empirical distribution ${\mathbb{P}}_{N}=\frac{1}{N}\sum_{i\in[N]}\delta_{\bm{\xi}^{i}}$ where $\{\bm{\xi}^{i}=(\bm{x}^{i},y^{i})\}_{i\in[N]}$ are i.i.d. samples from $\mathbb{P}^{0}$ and $\delta_{\bm{\xi}}$ denotes the Dirac distribution supported on $\bm{\xi}$. The empirical risk minimization (ERM) problem is thus given by
$\begin{array}[]{cl}\underset{\bm{\beta}\in\mathbb{R}^{n}}{\inf}\mathbb{E}_{{\mathbb{P}}_{N}}[\ell_{\bm{\beta}}(\bm{x},y)]\;=\;\underset{\bm{\beta}\in\mathbb{R}^{n}}{\inf}\frac{1}{N}\sum_{i\in[N]}\ell_{\bm{\beta}}(\bm{x}^{i},y^{i}).\end{array}$ (4)
Distributionally Robust Optimization
As summarized in the introduction, DRO is motivated by the fact that in the finite-data setting, the distance between the true and empirical distributions is upper-bounded by some $\epsilon>0$, that is, $\mathbb{P}^{0}\in\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})$. The goal in DRO is to optimize the expected loss over the worst possible realization of a distribution residing in $\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})$:
@@
$\begin{array}[]{cl}\displaystyle\underset{\bm{\beta}\in\mathbb{R}^{n}}{\inf}&\displaystyle\underset{\mathbb{Q}\in\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})}{\sup}\quad\mathbb{E}_{\mathbb{Q}}[\ell_{\bm{\beta}}(\bm{x},y)].\end{array}$ (6)
We refer to [40] and [34] for the generalization guarantees and ML applications of 6.
Adversarial Robustness
@@
The goal of adversarial robustness is to provide robustness against adversarial attacks [28]. An adversarial attack, in the widely studied $\ell_{p}$-noise setting [18], perturbs the features of the test instances $(\bm{x},y)$ by adding additive noise $\bm{z}$ to $\bm{x}$. The adversary chooses the noise vector $\bm{z}$, subject to $\lVert\bm{z}\rVert_{p}\leq\alpha$, so as to maximize the loss $\ell_{\bm{\beta}}(\bm{x}+\bm{z},y)$ associated with this perturbed test instance. Therefore, ARO solves the following optimization problem in the training stage to hedge against adversarial perturbations at the test stage:

### v2

@@ Abstract
Adversarially robust optimization (ARO) has emerged as the de facto standard for training models that hedge against adversarial attacks in the test stage. While these models are robust against adversarial attacks, they tend to suffer severely from overfitting. To address this issue, some successful methods replace the empirical distribution in the training stage with alternatives including (i) a worst-case distribution residing in an ambiguity set, resulting in a distributionally robust (DR) counterpart of ARO; (ii) a mixture of the empirical distribution with a distribution induced by an auxiliary (e.g., synthetic, external, out-of-domain) dataset. Inspired by the former, we study the Wasserstein DR counterpart of ARO for logistic regression and show it admits a tractable convex optimization reformulation. Adopting the latter setting, we revise the DR approach by intersecting its ambiguity set with another ambiguity set built using the auxiliary dataset, which offers a significant improvement whenever the Wasserstein distance between the data generating and auxiliary distributions can be estimated. We study the underlying optimization problem, develop efficient solution algorithms, and demonstrate that the proposed method outperforms benchmark approaches on standard datasets.
@@ 1 INTRODUCTION
Supervised learning traditionally involves access to a training dataset whose instances are assumed to be independently sampled from a true data-generating distribution (Bishop,, 2006; Hastie et al.,, 2009). Optimizing an expected loss for the empirical distribution constructed from such a training set, also known as empirical risk minimization (ERM), enjoys several desirable properties in relatively generic settings, including convergence to the true risk minimization problem as the number of training samples increases (Vapnik,, 1999, Chapter 2). In real-world applications, however, various challenges, such as data scarcity and the existence of adversarial attacks, lead to deteriorated out-of-sample performance for models trained via ERM.
One of the key limitations of ERM, particularly as it is designed to minimize an expected loss for the empirical distribution, emerges from the finite nature of data in practice. This leads ERM to suffer from the ‘optimism bias’, also known as overfitting (Murphy,, 2022), or the optimizer’s curse (DeMiguel and Nogales,, 2009; Smith and Winkler,, 2006), causing deteriorated out-of-sample performance. A popular approach to prevent this phenomenon, distributionally robust optimization (DRO; Delage and Ye, 2010), optimizes the expected loss for the worst-case distribution residing within a pre-specified ambiguity set.
Another key challenge faced by ERM in practice is adversarial attacks, where an adversary perturbs the observed features during the testing or deployment phase (Szegedy et al.,, 2014; Goodfellow et al.,, 2015). For neural networks, the paradigm of adversarial training (AT; Madry et al., 2018) is thus designed to provide adversarial robustness by simulating the attacks in the training stage. Several successful variants of AT, specialized to different losses and attacks, have been proposed in the literature to achieve adversarial robustness without significantly reducing performance on training sets (Shafahi et al.,, 2019; Zhang et al.,, 2019; Pang et al.,, 2022; Gao et al.,, 2019). While some works (e.g., Carlini et al., 2019; Uesato et al., 2018) examine adversarial robustness guarantees of various training algorithms, a recent stream of research (e.g., Bennouna et al., 2023; Xing et al., 2022b ) focuses on adversarially robust optimization (ARO), constraining ERM to guarantee an exact pre-specified level of adversarial robustness while maximizing training accuracy.
Recently, it has been observed that the two aforementioned notions of robustness can conflict, as adversarially robust (AR) models suffer from severe overfitting (robust overfitting; Raghunathan et al., 2019; Yu et al., 2022; Li and Spratling, 2023). Indeed, it is observed that robust overfitting is even more severe than traditional overfitting (Rice et al.,, 2020). To this end, some works address robust overfitting by revisiting AT algorithms and adding adjustments for better generalization (Chen et al.,, 2020; Li and Li,, 2023). In a recent work, Bennouna et al., (2023, Thm 3.2) decompose the error gap of robust overfitting into the statistical error of estimating the true data-generating distribution via the empirical distribution and an adversarial error resulting from the adversarial attacks, hence proposing the simultaneous adoption of DRO and ARO.
In this work, we study logistic regression (LR) for binary classification that is adversarially robust against $\ell_{p}$-attacks (Croce et al.,, 2020). To address robust overfitting faced by the adversarially robust LR model, we employ a DRO approach where distributional ambiguity is modeled with the type-1 Wasserstein metric. We base our work on an observation that the worst-case logistic loss under adversarial attacks can be represented as a Lipschitz continuous and convex loss function. This allows us to use existing Wasserstein DRO machinery for Lipschitz losses, and derive an exact reformulation of the Wasserstein DR counterpart of adversarially robust LR as a tractable convex problem.
Our main contribution lies in reducing the size of the Wasserstein ambiguity set in the DRO problem mentioned above, in order to create a less conservative problem while preserving the same robustness guarantees. To accomplish this, we draw inspiration from recent work on ARO that leverages auxiliary datasets (e.g., Gowal et al., 2021; Xing et al., 2022b ) and revise our DRO problem by intersecting its ambiguity set with another ambiguity set constructed using an auxiliary dataset. Examples of auxiliary data include synthetic data generated from a generative model (e.g., privacy-preserving data release), data in the presence of distributional shifts (e.g., different time periods/regions), noisy data (e.g., measurement errors), or out-of-domain data (e.g., different source); any auxiliary dataset is viable as long as its instances are sampled independently from an underlying data-generating distribution whose Wasserstein distance to the true data-generating distribution is known or can be estimated. Figure 1 illustrates our framework.
@@
Figure 1: Traditional ARO optimizes the expected adversarial loss over the empirical distribution $\mathbb{P}_{N}$ constructed from $N$ i.i.d. samples of the (unknown) true data-generating distribution $\mathbb{P}^{0}$. Replacing $\mathbb{P}_{N}$ with a worst-case distribution in the ball $\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})$ gives us its DR counterpart. To reduce the size of this ball, we intersect it with another ball $\mathfrak{B}_{\widehat{\varepsilon}}(\widehat{P}_{\widehat{N}})$ while ensuring $\mathbb{P}^{0}$ is still included with high confidence. The latter ball is centered at an empirical distribution ^$\widehat{\mathbb{P}}_{N}$ constructed from $\widehat{N}$ i.i.d. samples of some auxiliary distribution $\widehat{\mathbb{P}}$. Recent works using auxiliary data in ARO propose optimizing the expected adversarial loss over a mixture $\mathbb{Q}_{\mathrm{mix}}$ of $\mathbb{P}_{N}$ and $\widehat{P}_{\widehat{N}}$; we show that this distribution resides in $\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}({\widehat{\mathbb{P}}}_{\widehat{N}})$ under some conditions.
The paper unfolds as follows. In Section 2, we review related literature on DRO and ARO, with a focus on their interactions. We examine the use of auxiliary data in ARO and the intersection of Wasserstein balls in DRO. We discuss open questions for LR to motivate our loss function choice in this work. Section 3 gives preliminaries on ERM, ARO, and type-1 Wasserstein DRO. In Section 4, we discuss that the adversarial logistic loss can be reformulated as a Lipschitz convex function, enabling the use of Wasserstein DRO machinery for Lipschitz losses. Our main contribution (cf. Figure 1) is in Section 5, where we provide an explicit reformulation of the distributionally and adversarially robust LR problem over the intersection of two Wasserstein balls, prove the NP-hardness of this problem, and derive a convex relaxation of it. Our work is mainly on optimization where we focus on how to solve the underlying problems upon cross-validating Wasserstein ball radii, however, in Section 6 we discuss some preliminary statistical approaches to set such radii. We close the paper with numerical experiments on standard benchmark datasets in Section 7. We borrow the standard notation in DR machine learning, which is elaborated on in our Appendices.
@@ 2 RELATED WORK
Auxiliary data in ARO. The use of auxiliary data appears in the ARO literature. In particular, it is shown that additional unlabeled data sampled from the same (Carmon et al.,, 2019; Xing et al., 2022a, ) or different (Deng et al.,, 2021) data-generating distributions could provide adversarial robustness. Sehwag et al., (2022) show that adversarial robustness can be certified even when it is provided for a synthetic dataset as long as the distance between its generator and the true data-generating distribution can be quantified. Gowal et al., (2021); Xing et al., 2022b propose optimizing a weighted combination of ARO over empirical and synthetic datasets. We show that the latter approach can be recovered by our model.
@@
DRO-ARO interactions. In our work, we optimize ARO against worst-case data-generating distributions residing in an ambiguity set, where the type-1 Wasserstein metric is used for distances since it is arguably the most common choice in machine learning (ML) with Lipschitz losses (Shafieezadeh-Abadeh et al.,, 2019; Gao,, 2023). In the literature, it is shown that standard ARO is equivalent to the DRO of the original loss function with a type-$\infty$ Wasserstein metric (Staib and Jegelka,, 2017; Khim and Loh,, 2018; Pydi and Jog,, 2021; Regniez et al.,, 2022; Frank and Niles-Weed,, 2024), or a Lévy-Prokhorov metric (Bennouna and Van Parys,, 2022). In other words, in the absence of adversarial attacks, training models adversarially with artificial attacks provide some distributional robustness. Hence, our DR ARO approach can be interpreted as optimizing the logistic loss over the worst-case distribution whose $1$-Wasserstein distance is bounded by a pre-specified radius from at least one distribution residing in an $\infty$-Wasserstein ball around the empirical distribution. Conversely, Sinha et al., (2018) discuss that while DRO over Wasserstein balls is intractable for generic losses (e.g., neural networks), its Lagrange relaxation resembles ARO and thus ARO yields a certain degree of (relaxed) distributional robustness (Wu et al.,, 2020; Bui et al.,, 2022; Phan et al.,, 2023). However, to the best of our knowledge, there have not been works optimizing a pre-specified level of type-$1$ Wasserstein distributional robustness (that hedges against overfitting, Kuhn et al., 2019) and adversarial robustness (that hedges against adversarial attacks, Goodfellow et al., 2015) simultaneously. To our knowledge, the only approach that considers the exact DR counterpart of ARO is proposed by Bennouna et al., (2023) who model distributional ambiguity with $\varphi$-divergences for neural networks.
@@
Intersecting ambiguity sets in DRO. Recent work started to explore the intersection of ambiguity sets for different contexts (Awasthi et al.,, 2022; Wang et al.,, 2024) or different metrics (Zhang et al.,, 2023). Our idea of intersecting Wasserstein balls is originated from the “Surround, then Intersect” strategy (Taskesen et al.,, 2021, §5.2) to train linear regression under sequential domain adaptation in a non-adversarial setting (see Shafahi et al., 2020 and Song et al., 2019 for robustness in domain adaptation/transfer learning). The aforementioned work focuses on the squared loss function with an ambiguity set using the Wasserstein metric developed for the first and second distributional moments. In a recent study, Rychener et al., (2024) generalize most of the previous results and prove that DRO problems over the intersection of two Wasserstein balls admit tractable convex reformulations whenever the loss function is the maximum of concave functions
@@
Logistic loss in DRO and ARO. Our choice of LR aligns with the current directions and open questions in the related literature. In the DRO literature, even in the absence of adversarial attacks, the aforementioned work of Taskesen et al., (2021) on the intersection of Wasserstein ambiguity sets is restricted to linear regression. The authors show that this problem admits a tractable convex optimization reformulation, and their proof relies on the properties of the squared loss. Similarly, Rychener et al., (2024) discuss that the logistic loss fails to satisfy the piece-wise concavity assumption and is inherently difficult to optimize over the intersection of Wasserstein balls. We contribute to the DRO literature for adversarial and non-adversarial settings because we show that such a problem would be NP-hard for the logistic loss even without adversarial attacks, and develop specialized approximation techniques. Our problem recovers DR LR (Shafieezadeh-Abadeh et al.,, 2015; Selvi et al.,, 2022) as a special case in the absence of adversarial attacks and auxiliary data. Answering theoretical challenges posed by logistic regression has been useful in answering more general questions in the DRO literature, such as DR LR (Shafieezadeh-Abadeh et al.,, 2015) leading to DR ML (Shafieezadeh-Abadeh et al.,, 2019) and mixed-feature DR LR (Selvi et al.,, 2022) leading to mixed-feature DR Lipschitz ML (Belbasi et al.,, 2023). Finally, in the (non-DR) ARO literature, there are recent theory developments on understanding the effect of auxiliary data (e.g.Xing et al., 2022b ) specifically for squared and logistic loss functions.
3 PRELIMINARIES
@@
We consider a binary classification problem where an instance is modeled as $(\bm{x},y)\in\Xi:=\mathbb{R}^{n}\times\{-1,+1\}$ and the labels depend on the features via $\mathrm{Prob}[y\mid\bm{x}]\;=\;[1+\exp(-y\cdot\bm{\beta}^{\top}\bm{x})]^{-1}$ for some $\bm{\beta}\in\mathbb{R}^{n}$; its associated loss is the logloss $\ell_{\bm{\beta}}(\bm{x},y):=\log(1+\exp{(-y\cdot\bm{\beta}^{\top}\bm{x})})$.
@@
Empirical risk minimization. Let $\mathcal{P}(\Xi)$ denote the set of distributions supported on $\Xi$ and $\mathbb{P}^{0}\in\mathcal{P}(\Xi)$ denote the true data-generating distribution. One wants to minimize the expected logloss over $\mathbb{P}^{0}$, that is
$\begin{array}[]{cl}\displaystyle\underset{\bm{\beta}\in\mathbb{R}^{n}}{\inf}\;\displaystyle\mathbb{E}_{\mathbb{P}^{0}}[\ell_{\bm{\beta}}(\bm{x},y)].\end{array}$ (2)
@@
In practice, $\mathbb{P}^{0}$ is hardly ever known, and one resorts to the empirical distribution ${\mathbb{P}}_{N}=\frac{1}{N}\sum_{i\in[N]}\delta_{\bm{\xi}^{i}}$ where $\bm{\xi}^{i}=(\bm{x}^{i},y^{i}),\ i\in[N]$, are i.i.d. samples from $\mathbb{P}^{0}$ and $\delta_{\bm{\xi}}$ denotes the Dirac distribution supported on $\bm{\xi}$. The empirical risk minimization (ERM) problem is thus
$\begin{array}[]{cl}\displaystyle\underset{\bm{\beta}\in\mathbb{R}^{n}}{\inf}\;\mathbb{E}_{{\mathbb{P}}_{N}}[\ell_{\bm{\beta}}(\bm{x},y)].\end{array}$ (4)
Distributionally robust optimization. To be able to define a distance between distributions, we first define the following feature-label metric on $\Xi$.
Definition 1.
The distance between instances $\bm{\xi}=(\bm{x},y)\in\Xi$ and $\bm{\xi^{\prime}}=(\bm{x^{\prime}},y^{\prime})\in\Xi$ for $\kappa>0$ and $q\geq 1$ is
$d(\bm{\xi},\bm{\xi^{\prime}})=\lVert\bm{x}-\bm{x^{\prime}}\rVert_{q}+\kappa\cdot\mathds{1}{[y\neq y^{\prime}]}.$
Using this metric, we define the Wasserstein distance.
Definition 2.
The type-1 Wasserstein distance between distributions $\mathbb{Q},\mathbb{Q}^{\prime}\in\mathcal{P}(\Xi)$ is defined as
$\mathrm{W}(\mathbb{Q},\mathbb{Q}^{\prime})=\underset{\Pi\in\mathcal{C}(\mathbb{Q},\mathbb{Q}^{\prime})}{\inf}\left\{\int_{\Xi\times\Xi}d(\bm{\xi},\bm{\xi^{\prime}})\Pi(\mathop{}\!\mathrm{d}\bm{\xi},\mathop{}\!\mathrm{d}\bm{\xi^{\prime}})\right\},$
where $\mathcal{C}(\mathbb{Q},\mathbb{Q}^{\prime})$ is the set of couplings of $\mathbb{Q}$ and $\mathbb{Q}^{\prime}$.
In finite-data settings, the distance between the true data-generating distribution and the empirical distribution is upper-bounded by some $\epsilon>0$. The Wasserstein DRO problem is thus defined as
@@
$\begin{array}[]{cl}\displaystyle\underset{\bm{\beta}\in\mathbb{R}^{n}}{\inf}\;\displaystyle\underset{\mathbb{Q}\in\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})}{\sup}\;\mathbb{E}_{\mathbb{Q}}[\ell_{\bm{\beta}}(\bm{x},y)],\end{array}$ (6)
where $\mathfrak{B}_{\varepsilon}(\mathbb{P}):=\{\mathbb{Q}\in\mathcal{P}(\Xi):\mathrm{W}(\mathbb{Q},\mathbb{P})\leq\varepsilon\}$ denotes the Wasserstein ball centered at $\mathbb{P}\in\mathcal{P}(\Xi)$ with radius $\varepsilon$. We refer to Mohajerin Esfahani and Kuhn, (2018) and Kuhn et al., (2019) for the properties of 6.
Adversarially robust optimization.
@@
The goal of adversarial robustness is to provide robustness against adversarial attacks (Goodfellow et al.,, 2015). An adversarial attack, in the widely studied $\ell_{p}$-noise setting (Croce et al.,, 2020), perturbs the features of the test instances $(\bm{x},y)$ by adding additive noise $\bm{z}$ to $\bm{x}$. The adversary chooses the noise vector $\bm{z}$, subject to $\lVert\bm{z}\rVert_{p}\leq\alpha$, so as to maximize the loss $\ell_{\bm{\beta}}(\bm{x}+\bm{z},y)$ associated with this perturbed test instance. Therefore, ARO solves the following optimization problem in the training stage to hedge against adversarial perturbations at the test stage: